For the function y = x2 - 4x - 5, perform the following tasks:
a) Put the function in the form y = a(x - h)2 + k.
Answer:
Show work in this space



b) What is the equation for the line of symmetry for the graph of this function?
Answer:



c) Graph the function using the equation in part a. Explain why it is not necessary to plot points to graph when using y = a (x – h)2 + k.
Show graph here.


Explanation of graphing.




d) In your own words, describe how this graph compares to the graph of y = x2?
Answer:



Answer by vanetiks(6)

For the function y = x2 - 4x - 5, perform the following tasks:
a) Put the function in the form y = a(x - h)2 + k.
Answer: a=1, h=2, k=-9
Show work in this space
={x2-2(x)(2)+22}-22-5
=(x-2)2-9
b) What is the equation for the line of symmetry for the graph of this function?
Answer: x-2=0....or...x=2 is the line of symmetry
c) Graph the function using the equation in part a. Explain why it is not necessary to plot points to graph when using y = a (x – h)2 + k.
Show graph here.
I am sorry, but my graph won't show. I hope the rest of this info helps you. If you want to graph it yourself the first parabola curves upward touching the (0,0) like a U shape and the next parabola touches the points on the x axis at -1 and -5 and the vetex is at -9. It still curves upward like a U.
Explanation of graphing.
Draw line of symmetry x=2. Plot vertex at (2,-9). Plot curve symmetrically along the line of symmetry taking the 2 intercept points on the x axis as
x-2=+3 or -3....that is x=5 and -1
d) In your own words, describe how this graph compares to the graph of y = x2?
Answer: There is a shift in vertex from (0,0) to (2,-9)